OpenStudy (brittany):

Translate this problem to an equation and solve. The product of 2 positive consecutive even integers is 48. What is the smaller number

OpenStudy (anonymous):

Let n be the smaller integer. What will be the other one?

OpenStudy (brittany):

n+2?

OpenStudy (anonymous):

n+1 is the next integer.

OpenStudy (brittany):

oh ok

OpenStudy (anonymous):

n+2 would be the next odd/even integer but they don't ask about that.

OpenStudy (brittany):

n(n+1)=48

OpenStudy (anonymous):

Yeah, that last one. Because they said it's the product of those. Therefore you must use multiplication.

OpenStudy (brittany):

what do i do next?

OpenStudy (brittany):

2n+1=48?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

n(n+1) = 48

OpenStudy (brittany):

n^2+1=48?

OpenStudy (anonymous):

\[\implies n^2 + n =48\]

OpenStudy (brittany):

2n^2=48?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

You can't combine an n and an n^2 term

OpenStudy (brittany):

oh ok

OpenStudy (anonymous):

\[\implies n^2 + n - 48 = 0\] Then use the quadratic formula.

OpenStudy (brittany):

whats tht?

OpenStudy (anonymous):

Although this makes no sense since there are no integer solutions to this equation.

OpenStudy (anonymous):

oh crap sorry

OpenStudy (anonymous):

They did say 'even' integers.

OpenStudy (anonymous):

So it's n(n+2) = 48

OpenStudy (brittany):

oh hahaha

OpenStudy (anonymous):

\[\implies n^2 + 2n = 48\]\[\implies n^2 + 2n - 48 = 0\]

OpenStudy (brittany):

2n+2=48 -2 -2 2n+46 /2 /2

OpenStudy (anonymous):

Then you can factor that?

OpenStudy (anonymous):

no. You keep adding when you need to multiply.

OpenStudy (brittany):

Ugh what

OpenStudy (anonymous):

n(n+2) = n(n) + n(2)

OpenStudy (anonymous):

= n^2 + 2n

OpenStudy (anonymous):

Do you use twiddla at all?

OpenStudy (brittany):

twiddla?

OpenStudy (anonymous):

http://www.twiddla.com/591991

OpenStudy (anonymous):

it's a whiteboard so I can demonstrate better.

OpenStudy (brittany):

oh oksure?

OpenStudy (brittany):

hello

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

So \[n^2 + 2n - 48 = 0\]

OpenStudy (brittany):

yes

OpenStudy (anonymous):

So can you factor that from there?

OpenStudy (brittany):

factor?

OpenStudy (anonymous):

Like (n+a)(n+b) Where a + b = 2 And a*b = -48

OpenStudy (brittany):

i dont understand

OpenStudy (anonymous):

Factoring. Or solving quadratic equations

OpenStudy (anonymous):

Have you done this before?

OpenStudy (brittany):

no?

OpenStudy (anonymous):

Hrm. Well I guess you can solve this using some level of intuition.

OpenStudy (anonymous):

What two even numbers can you multiply to get 48 that are separated by a distance of 2?

OpenStudy (brittany):

8 and 6?

OpenStudy (anonymous):

Correct. And the smallest of those is?

OpenStudy (brittany):

6. oh my god thank you so much ahaha i apprectiat it

OpenStudy (anonymous):

No problem ;)